OpenStudy (anonymous):

Which expression is equivalent to one over fivem − 20? one over five(m − 4) one over five(m − 100) 5(m − 4) 5(m − 100) will give 2 medals

OpenStudy (anonymous):

Question 2 (Multiple Choice Worth 4 points) Write an expression that is equivalent to three over five (3y + 15). three over fivey + 9 three over five y + 15 nine over fivey + 9 nine over fivey + 15 Question 3 (Multiple Choice Worth 4 points) Write an expression for the perimeter of the triangle shown below: A triangle is shown with side lengths labeled 0.3x plus 24, 5x, and 0.2x minus 15. 10x − 9 10x − 39 5.5x − 9 5.5x + 9 Question 4 (Multiple Choice Worth 4 points) Monica purchased adult and youth tickets for the fair. She bought x adult tickets for $28 each and y youth tickets for $15. Write an expression that can be used to show the total cost, C, of all tickets. C = (28 + 15)x C = (28 + 15)(x + y) C = 28x + 15y C = 28y + 15x Question 5 (Multiple Choice Worth 4 points) Ella purchased a game that was on sale for 12% off. The sales tax in her county is 6%. Let y represent the original price of the game. Write an expression that can be used to determine the final cost of the game. y − 0.12y 0.06(0.88y) 1.06(0.88y) y − 0.88y + 0.06y